# Text2VDM: Text to Vector Displacement Maps for Expressive and Interactive 3D Sculpting

### v1

@@ Abstract
Professional 3D asset creation often requires diverse sculpting brushes to add surface details and geometric structures. Despite recent progress in 3D generation, producing reusable sculpting brushes compatible with artists’ workflows remains an open and challenging problem. These sculpting brushes are typically represented as vector displacement maps (VDMs), which existing models cannot easily generate compared to natural images. This paper presents Text2VDM, a novel framework for text-to-VDM brush generation through the deformation of a dense planar mesh guided by score distillation sampling (SDS). The original SDS loss is designed for generating full objects and struggles with generating desirable sub-object structures from scratch in brush generation. We refer to this issue as semantic coupling, which we address by introducing classifier-free guidance (CFG) weighted blending of prompt tokens to SDS, resulting in a more accurate target distribution and semantic guidance. Experiments demonstrate that Text2VDM can generate diverse, high-quality VDM brushes for sculpting surface details and geometric structures. Our generated brushes can be seamlessly integrated into mainstream modeling software, enabling various applications such as mesh stylization and real-time interactive modeling.
Figure 1: Example VDM brushes generated by Tex2VDM and sculpted models in Blender. Tex2VDM can produce high-quality brushes for surface details (top row) and geometric structures (bottom row) from text input. Users can rapidly create an expressive model from a plain shape by directly applying these brushes in Blender. Yellow underlined text highlights semantics enhanced by our framework.
@@ 1 Introduction
Despite significant advances in text-to-image (T2I) [40, 33] and text-to-3D generation [49, 20, 27, 48, 34], existing methods are unsuitable for creating VDM brushes. We summarize the challenges as follows: 1) Since VDMs are not natural images (Figure 2), it is difficult for existing T2I models to generate them directly. 2) From a 3D perspective, a VDM represents mesh deformation through per-vertex displacement vectors from a dense planar mesh. Mapping any generated mesh to a dense planar mesh to create a VDM is non-trivial. 3) Sculpting brushes often involve sub-object structures, whereas most 3D generation methods can only generate full objects. Enabling users to accurately control the generation of sculpting brush through text descriptions in a semantically focused manner remains challenging.
To address the challenges of brush generation, we propose Text2VDM, a novel optimization-based framework that generates diverse and controllable VDM brushes from text input. Our approach does not generate VDMs directly from a T2I model. Instead, we address VDM brush generation from a 3D perspective by applying score distillation with a pre-trained T2I model to guide mesh deformation. Our framework supports three ways to initialize a base mesh through a zero-valued, spike-pattern, or user-specified VDM for custom shape control. We reparameterize the mesh vertices through an implicit formulation based on the Laplace-Beltrami operator [32] to achieve high-quality optimization of mesh deformations. We also provide optional region control using a mask of activated mesh deformation, helping users obtain the intended brush effects. We then rasterize normal maps of the mesh using a differentiable renderer for brush optimization.
We observed that the standard score distillation sampling (SDS) [34] can lead to semantic coupling when supervising the generation of sub-object level structures because of associated semantics noisy gradients of full object. For example, a generated tortoiseshell should not be a full tortoise with a head and a tail. A straightforward solution is to use negative prompts [57, 17] to exclude undesired semantics, but our experiments show that this approach is ineffective in decoupling semantics and leads to an unstable and more time-consuming optimization process. Instead, we propose to enhance the semantics of part-related words by applying classifier-free guidance (CFG) weighted blending to the tokens in the prompt. This results in semantically focused text embedding, directing toward a more precise target distribution while reducing noisy gradients during optimization.
@@
Our experiments demonstrate that Text2VDM produces high-quality and diverse VDM brushes that can be directly integrated into mainstream modeling software, such as Blender [5] and ZBrush [11]. Compared to existing methods that directly generate full 3D models, our approach addresses a different use case where brush-based user sculpting is desirable. We enable artists to interactively use a variety of brushes to sculpt diverse and expressive models from a plain shape.
@@
  • We introduce CFG-weighted blending to SDS for modeling a more precise target distribution, mitigating semantic coupling in sub-object structure generation.
Figure 3: Overview of Text2VDM. Starting with a dense planar mesh constructed from a zero-valued VDM, users can initialize the mesh volume through a default spike-pattern VDM or a user-specified VDM. Given the text prompt and region mask, we apply CFG-weighted SDS loss $\mathcal{L}_{\text{WS}}$ to guide mesh deformations through a Laplace-Beltrami operator $L$ iteratively, achieving semantically focused generation of surface details or geometric structures. After optimization, vertex displacements are baked into the final VDM.
@@ 2 Related Work
This section reviews previous work related to 3D sculpting brush generation and summarizes the current research gap.
2.1 Text to Local 3D Generation and Editing
@@
With recent advances in diffusion models [39] and differentiable 3D representations [43, 45, 1, 30, 32], many methods for text-guided full 3D model generation have emerged [7, 10, 20, 35, 22, 41, 28]. Since 3D content creation is an iterative process that often requires user interaction, more attention has been directed toward localized 3D generation and editing. For example, 3D Highlighter [8] and 3D Paintbrush [9] use text as input, leveraging pre-trained CLIP models [38] or diffusion models [34] to supervise the optimization of neural networks for segmenting the regions of a 3D model that match the text description. Based on the information from these segmented regions, further editing of texture and geometry can be applied to the 3D model. Furthermore, SKED [29] and SketchDream [24] introduce sketches as an additional modality to assist in localized editing. To enable more precise control, FocalDreamer [21], MagicClay [4], and Tip-Editor [58] allow users to specify the editing location directly within the 3D space. These works rely on optimization-based methods to edit specific objects, often resulting in non-reusable editing outcomes. Additionally, each edit requires a lengthy optimization process, making interactivity difficult to achieve.
2.2 Diffusion Priors for 3D Generation
@@
Score distillation sampling (SDS) [51, 34] provides pixel-level guidance by seeking specific modes in a diffusion model, inspiring further research to improve optimization-based 3D generation [53, 23, 3, 52, 55]. Some studies focus on mitigating the “Janus” problem [2, 14], while others fine-tune diffusion models with multi-view datasets to enhance 3D consistency [44, 25]. Recent research focuses on refining the design of SDS loss to achieve more precise guidance. For instance, Make-it-3D [49] introduces two-stage optimizations to improve textured appearance, while Fantasia3D [7] dynamically modifies the time-dependent weighting function within SDS computations. Additionally, several methods [57, 17] incorporate negative prompts as the conditional term to further refine the optimizations. Although diffusion priors have achieved promising results, their application in generating sub-object structures without global context as a reference is still challenging.
2.3 Appearance and Geometric Brush Synthesis
@@
The concept of brushes is very common in the creative process of digital artists, serving as a reusable local decorative unit. Appearance brushes focus on color representation and drawing styles in 2D space. With the development of generative models [12, 39], many works have explored the synthesis appearance brushes for interactive painting [15, 46], realistic artworks generation [59, 26, 31], and applying stylization [18, 16]. Unlike appearance brushes, geometric brushes focus on modifying geometry by moving the vertices of a mesh in 3D space. VDM brushes, as an extension of standard geometric brushes, provide more complex geometric effects by utilizing VDMs. To the best of our knowledge, only a few techniques adopted the concepts of VDM for generation [54, 42] and geometric texture transfer [13]. Generating geometric brushes that can be used within existing workflows is still under-explored.
@@ 3 Methodology
Our framework can generate VDM brushes compatible with mainstream modeling software. As shown in Figure 3, we begin by constructing a dense mesh $\mathcal{M}_{0}$ from an initial VDM $T_{0}$. We then apply score distillation with a pretrained T2I diffusion model to guide mesh deformation through a Laplace-Beltrami operator $L$ [32] toward the input text prompt $y$. The deformable region of the mesh can be controlled by an optional user-specified region mask $R$. To produce a high-quality sub-object level structure that only represents the intended part described by the text, we apply CFG-weighted blending to the tokens in the prompt, effectively handling the issue of semantic coupling in SDS.
@@ 3.1 Brush Initialization
We provide three methods to initialize a base mesh for brush generation via a zero-valued VDM, a spike-pattern VDM, or a user-specified VDM. A VDM is represented as a $512\times 512$ three-channel image, in which each channel stores the displacement in the X, Y, or Z direction respectively. We first construct a planar grid mesh by creating two triangles for every $2\times 2$ pixels and then apply the displacement stored in the VDM to mesh vertices. The values in these three initial VDMs range from 0 to 1, in which 0 represents no displacement, and 1 corresponds to half of the mesh’s edge length in the positive axis direction. Since users can apply sculpting brushes symmetrically, our initial VDM does not need to store any negative values.
@@ 3.2 Brush Generation via Mesh Deformation
Given the initialized based mesh, our method aims to learn a mesh deformation to the target brush shape. The vertex positions $\hat{v}$ after mesh deformation can be expressed by
$\hat{v}=\operatorname*{arg\,min}_{v}\mathcal{L}_{\text{WS}}(\mathcal{D}_{c}(v),y),$ (1)
where $c$ represents the camera setup in a differentiable renderer $\mathcal{D}$ [19]. The loss function $\mathcal{L}_{\text{WS}}$ receives the rendered normal image $\mathcal{D}_{c}(v)$ and text input $y$ to evaluate the semantic guidance, which is detailed in Section 3.3. In mesh deformation strategies, directly applying displacement to each vertex often results in unintended self-intersections of mesh faces caused by noisy gradients from pixel-level losses [28]. To address it, several works [10, 50] adopt the strategy by Aigerman et al. [1], parameterizing deformation by Jacobian fields that capture the scaling and rotation of each mesh face. Although this method effectively smooths vertex displacements, the local deformation represented in Jacobians accumulates, leading to global drifting for open-boundary meshes, making it challenging to bake the mesh as a brush.
Based on these observations, we reparameterize the vertex optimization in Equation 1 through an implicit formulation with a Laplace-Beltrami operator $L$, similar to the approaches presented by Nicolet et al. [32]:
$v^{*}=(I+\lambda L)v,$ (2)
where $I$ is the identity matrix, $\lambda$ is a hyperparameter to control the extent of gradient diffusion from a given vertex to its neighboring vertices. When $\lambda=0$, this representation degrades to direct vertex displacements. As $\lambda$ increases, the mesh deforms toward more global structural changes. In our experiments, we set $\lambda=15$ to balance the global structure and fine details during mesh deformation (Figure 4). Additionally, we provide a region mask to restrict mesh deformation to the user-defined region during optimization. By adjusting the activation ratio of the region mask, the final brush effect can effectively match the user’s guidance. For instance, our experiment activated the region mask for the first half of total iterations as a warm-up stage to effectively control the shape of the surface detail (Figure 8). This differentiable parameterization effectively alters the gradient propagation at each optimization step as:
$v^{*}\leftarrow v^{*}-\eta(I-\lambda L)^{-1}\frac{\partial{\mathcal{L}_{\text{WS}}}}{\partial{v}^{*}}$ (3)
where $\eta$ is the learning rate. The advantage of this Laplacian energy-aware mesh deformation is that it enhances the robustness of optimization for non-convex objective functions. The resulting mesh preserves the original structure while incorporating rich local deformations, making it well-suited for baking as a brush.
Figure 4: Effects of $\lambda$. As $\lambda$ decreases, the mesh deformation shifts from global structural to locality, resulting in finer details, until it becomes a noisy surface at $\lambda$ = 0.
3.3 CFG-Weighted Score Distillation Sampling
@@
Current text-to-3D generation methods like DreamFusion [34] often optimize a 3D representation parameterized by $\theta$ so that rendered images $\mathbf{x}=g(\theta)$ resemble 2D samples produced by a pre-trained T2I diffusion model for a given text prompt $y$. $g$ functions as a differentiable renderer. The T2I diffusion model $\phi$ predicts the sampled noise $\epsilon_{\phi}(\mathbf{x}_{t};y,t)$ of a rendered image $\mathbf{x}_{t}$ at a noise level $t$ for the text input $y$. To move all rendered images with higher density regions under the text-conditioned diffusion prior, SDS loss estimates the gradient for updating $\theta$ as:
@@
However, directly supervising sub-object structures with SDS, can lead to semantic coupling, which results in the generation of extra semantically related parts. For example, when generating a tortoiseshell, this semantic coupling will cause the generation of the tortoise’s tail and head (Figure 7). We believe that the issue of semantic coupling in SDS stems from the training data of the stable diffusion model, where images often depict partial components along with the complete object. This causes the target distribution conditioned by the text description of sub-object structures containing semantic information related to the full object.
@@
A straightforward approach is using negative prompts proposed by classifier score distillation (CSD) [57] to mitigate coupled semantics. CSD demonstrates that compared to variational score distillation (VSD) [53], which adaptively learns negative classifier scores, CSD employs predefined negative prompts, resulting in a more precise optimization process:
@@
$\nabla_{\theta}\mathcal{L}_{\text{CSD}}(\phi,\mathbf{x})$ $=\mathbb{E}_{t,\epsilon,c}[(\omega_{1}\cdot\epsilon_{\phi}(\mathbf{x}_{t};y,t)$
@@
$-\omega_{2}\cdot\epsilon_{\phi}(\mathbf{x}_{t};y_{neg},t))\frac{\partial\mathbf{x}}{\partial\theta}],$ (5)
where $\omega_{1}$ and $\omega_{2}$ denote different weights for positive and negative prompts. We found that the negative prompt’s semantic distribution does not align with the associated semantics in the target distribution conditioned by the positive prompt of sub-object structures. This resulted in noisier gradients, making CSD less effective at decoupling semantics. Furthermore, as the weight of the negative prompt increased, the optimization became more unstable and challenging to converge.
Unlike CSD, we apply CFG-weighted blending to the tokens in the original prompt, which does not require additional inference to construct a negative distribution. This results in semantically focused text embedding, directing toward a more precise target distribution. Specifically, our loss function is defined as:
@@
$\small\nabla_{\theta}\mathcal{L}_{\text{WS}}(\phi,\mathbf{x})=\mathbb{E}_{t,\epsilon,c}\left[\omega(t)\left(\epsilon_{\phi}(\mathbf{x}_{t};y^{*},t)-\epsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right],$ (6)
@@
where $y^{*}$ is a text embedding computed by Compel [47]. Specifically, we assign each word in the prompt a CFG weight $s$ and compute the weighted embedding $e_{w}$ for each word by blending the original text embedding $e$ with the empty text embedding $e_{\phi}$ as follows: $e_{w}=e_{\phi}+s\cdot(e-e_{\phi})$. By concatenating the weighted embeddings of each word in sequence, we obtain the final semantically focused text embedding $y^{*}$. In our experiments, we found that assigning a weight of 1.21 to words that require enhanced semantics can achieve stable optimizations and effectively alleviate the issue of semantic coupling. Notably, the CFG weights used for text embedding computation are separate from the CFG guidance scale applied during the computation of the SDS loss. Our experiment uses a CFG guidance scale of 100.
@@
Figure 5: Qualitative comparisons of generated brushes for surface details. Our method captures geometry details guided by texts, effectively preserving surface structure and avoiding mesh distortions.

### v2

@@ Abstract
Professional 3D asset creation often requires diverse sculpting brushes to add surface details and geometric structures. Despite recent progress in 3D generation, producing reusable sculpting brushes compatible with artists’ workflows remains an open and challenging problem. These sculpting brushes are typically represented as vector displacement maps (VDMs), which existing models cannot easily generate compared to natural images. This paper presents Text2VDM, a novel framework for text-to-VDM brush generation through the deformation of a dense planar mesh guided by score distillation sampling (SDS). The original SDS loss is designed for generating full objects and struggles with generating desirable sub-object structures from scratch in brush generation. We refer to this issue as semantic coupling, which we address by introducing weighted blending of prompt tokens to SDS, resulting in a more accurate target distribution and semantic guidance. Experiments demonstrate that Text2VDM can generate diverse, high-quality VDM brushes for sculpting surface details and geometric structures. Our generated brushes can be seamlessly integrated into mainstream modeling software, enabling various applications such as mesh stylization and real-time interactive modeling.
Figure 1: Example VDM brushes generated by Text2VDM and sculpted models in Blender. Text2VDM can produce high-quality brushes for surface details (top row) and geometric structures (bottom row) from text input. Users can rapidly create an expressive model from a plain shape by directly applying these brushes in Blender. Yellow underlined text highlights semantics enhanced by our framework.
@@ 1 Introduction
Despite significant advances in text-to-image (T2I) [41, 36] and text-to-3D generation [50, 23, 30, 49, 37], existing methods are unsuitable for creating VDM brushes. We summarize the challenges as follows: 1) Since VDMs are not natural images (Figure 2), it is difficult for existing T2I models to generate them directly. 2) From a 3D perspective, a VDM represents mesh deformation through per-vertex displacement vectors from a dense planar mesh. Mapping any generated mesh to a dense planar mesh to create a VDM is non-trivial. 3) Sculpting brushes often involve sub-object structures, whereas most 3D generation methods can only generate full objects. Enabling users to accurately control the generation of sculpting brushes through text prompts in a semantically focused manner remains challenging.
To address the challenges of brush generation, we propose Text2VDM, a novel optimization-based framework that generates diverse and controllable VDM brushes from text input. Our approach does not generate VDMs directly from a T2I model. Instead, we address VDM brush generation from a 3D perspective by applying score distillation with a pre-trained T2I model to guide mesh deformation. Our framework supports three ways to initialize a base mesh through a zero-valued, spike-pattern, or user-specified VDM for custom shape control. For mesh deformation, we formulate a Sobolev preconditioned optimization [35] to maintain mesh quality with intrinsic smoothness. We also provide optional region control using a mask of activated mesh deformation, helping users obtain the intended brush effects. The normal maps of the mesh are then rasterized by a differentiable renderer for brush optimization.
We observed that the standard score distillation sampling (SDS) [37] can lead to semantic coupling when supervising the generation of sub-object level structures due to the associated semantics caused by the noisy gradients from the full object. For example, a generated deer’s horn should not be a full deer’s head, or a generated beard should not include a nose. A straightforward solution is to use negative prompts [59, 18] to exclude undesired semantics, but our experiments show that this semantics suppression approach is ineffective in decoupling semantics and leads to an unstable optimization process. Instead, we propose to enhance the semantics of part-related words by applying weighted blending to the tokens in the prompt. This results in semantically focused text embedding, directing toward a more precise target distribution while reducing noisy gradients during optimization.
@@
Our experiments demonstrate that Text2VDM produces high-quality and diverse VDM brushes that can be directly integrated into mainstream modeling software, such as Blender [5] and ZBrush [12]. Compared to existing methods that directly generate full 3D models, our approach addresses a different use case where brush-based user sculpting is desirable. This enables artists to interactively use a variety of brushes to sculpt diverse and expressive models from a plain shape.
@@
  • We design a novel Semantic Enhancement SDS loss, which uses weighted blending to mitigate semantic coupling for sub-object structure generation.
Figure 3: Overview of Text2VDM. Starting with a dense planar mesh constructed from a zero-valued VDM, users can initialize the mesh volume through a default spike-pattern VDM or a user-specified VDM. Given the text prompt and region mask, we propose the Semantic Enhancement SDS loss $\mathcal{L}_{\text{SE}}$ to guide mesh deformations through a mesh Laplacian $L$ iteratively, achieving semantically focused generation of surface details or geometric structures. After generation, vertex displacements are baked into the final VDM.
@@ 2 Related Work
Text to Local 3D Generation and Editing. With recent advances in diffusion models [41] and differentiable 3D representations [44, 46, 1, 33, 35], many methods for text-guided full 3D model generation have emerged [8, 11, 23, 38, 25, 42, 31]. Since 3D content creation is an iterative process that often requires user interaction, more attention has been directed toward localized 3D generation and editing. For example, 3D Highlighter [9] and 3D Paintbrush [10] use text as input, leveraging pre-trained CLIP models [40] or diffusion models [37] to supervise the optimization of neural networks for segmenting the regions of a 3D model that match the text description. Based on the information from these segmented regions, further editing of texture and geometry can be applied to the 3D model. Furthermore, SKED [32] and SketchDream [27] introduce sketches as an additional modality to assist in localized editing. To enable more precise control, FocalDreamer [24], MagicClay [4], and Tip-Editor [60] allow users to specify the editing location directly within the 3D space. These works rely on optimization-based methods to edit specific objects, often resulting in non-reusable editing outcomes. Additionally, each edit requires a lengthy optimization process, making interactivity difficult to achieve.
@@
Diffusion Priors for 3D Generation. Score distillation sampling (SDS) [52, 37] provides pixel-level guidance by seeking specific modes in a diffusion model, inspiring further research to improve optimization-based 3D generation [54, 26, 3, 53, 56]. Some studies focus on mitigating the “Janus” problem [2, 15], while others fine-tune diffusion models with multiview datasets to enhance 3D consistency [45, 28]. Recent research focuses on refining the design of SDS loss to achieve more precise guidance. For instance, Make-it-3D [50] introduces two-stage optimizations to improve textured appearance, while Fantasia3D [8] dynamically modifies the time-dependent weighting function within SDS computations. Additionally, several methods [59, 18] incorporate negative prompts as the conditional term to further refine the optimizations. Although diffusion priors have achieved promising results, their application in generating sub-object structures without global context as a reference is still challenging.
@@
Appearance and Geometric Brush Synthesis. The concept of brushes is very common in the creative process of digital artists, serving as a reusable local decorative unit. Appearance brushes focus on color representation and drawing styles in 2D space. With the development of generative models [13, 41], many works have explored the synthesis of procedural material [21, 22] for 3D object texturing and appearance brushes for interactive painting [16, 47], realistic artworks generation [61, 29, 34], and applying stylization [19, 17]. Unlike appearance brushes, geometric brushes focus on modifying geometry by moving the vertices of a mesh in 3D space. VDM brushes, as an extension of standard geometric brushes, provide more complex geometric effects by utilizing VDMs. To the best of our knowledge, only a few techniques adopted the concepts of VDM for generation [55, 43] and geometric texture transfer [14]. Recently, concurrent work [57] explored using a single image as input and leveraged a diffusion model for generating multiview normal maps to guide VDM reconstruction. In comparison, our method is fundamentally different in tackling the semantic coupling problem arising from text guidance. This demonstrates that generating geometric brushes is a highly promising research direction.
@@ 3 Methodology
To generate VDM brushes compatible with mainstream modeling software, we begin by constructing a dense mesh from an initial VDM, as shown in Figure 3. We then apply score distillation with Stable Diffusion to guide high-quality mesh deformation formulated as a form of Sobolev preconditioned optimization [35]. To produce the intended sub-object level structure described in the text, we design a Semantic Enhancement SDS loss by applying weighted blending to the tokens in the prompt, effectively handling the issue of semantic coupling in SDS.
@@ 3.1 Brush Initialization
We provide three methods to initialize a base mesh for brush generation via a zero-valued VDM, a spike-pattern VDM, or a user-specified VDM. A VDM is represented as a $512\times 512$ three-channel image, in which each channel stores the displacement in the X, Y, or Z direction, respectively. We first construct a planar grid mesh by creating two triangles for every $2\times 2$ pixels and then apply the displacement stored in the VDM to mesh vertices. The values in these three initial VDMs range from 0 to 1, in which 0 represents no displacement, and 1 corresponds to half of the mesh’s edge length in the positive axis direction. Since users can apply sculpting brushes symmetrically, our initial VDM does not need to store any negative values.
@@ 3.2 Brush Generation via Mesh Deformation
Given the vertices $v$ on the initialized base mesh, our method aims to learn a mesh deformation to the target brush shape. The vertex positions $\hat{v}$ after mesh deformation can be expressed by:
$\hat{v}=\operatorname*{arg\,min}_{v}\mathcal{L}_{\text{SE}}(\mathcal{D}_{c}(v),y),$ (1)
where $c$ represents the camera setup in a differentiable renderer $\mathcal{D}$ [20]. The loss function $\mathcal{L}_{\text{SE}}$ receives the rendered normal image $\mathcal{D}_{c}(v)$ and text input $y$ to evaluate the semantic guidance, which is detailed in Section 3.3. To accompany the external forces from $\mathcal{L}_{\text{SE}}$ that drive the mesh deformation with intrinsic smoothness energies, we follow the framework of a Sobolev preconditioned gradient descent [35], where the base mesh is reparameterized by the mesh Laplacian $L$:
$v^{*}=(I+\lambda L)v.$ (2)
This preconditioning involves solving a sparse linear system at every iteration, modifying the gradient descent update for each mesh deformation step to:
$v\leftarrow v-\eta(I+\lambda L)^{-1}\frac{\partial{\mathcal{L}_{\text{SE}}}}{\partial{v}},$ (3)
where $\eta$ is the learning rate, $I$ is the identity matrix, and $\lambda$ is a hyperparameter to control the extent of gradient diffusion over the entire domain. We set $\lambda=15$ throughout our experiments to balance the global structure and fine details during mesh deformation.
Compared to directly adding a Laplacian regularization term to Equation 1, the preconditioning framework [35] is critical in achieving large mesh deformation while maintaining proper topology with reduced triangle flips (Figure 4). Several works [11, 51] adopt the strategy by Aigerman et al. [1], parameterizing deformation through Jacobian fields that capture the scaling and rotation of each triangle. Although this method effectively smooths vertex displacements, the local deformation represented in Jacobians accumulates, leading to global drifting for open-boundary meshes, making it challenging to bake the mesh as a brush.
Additionally, we provide an optional region mask to restrict mesh deformation to the user-defined region during optimization. It helps maintain zero values in unused VDM areas when generating geometric structures. For producing surface details, the region mask ensures that the brush effect meets user-customized requirements (Figure 9).
By framing the VDM generation task as mesh deformation through preconditioned optimization, our framework preserves the favorable initial mesh topology and ensures control throughout the procedure. The resulting mesh preserves the original structure while incorporating rich local deformations, making it well-suited for baking as a brush.
Figure 4: Effect of $\lambda$. (a) Low-quality results w/o mesh smoothness, (b) low-quality results w/ direct Laplacian regularization, (c) our adopted preconditioning scheme with $\lambda$ = 15, and (d) results with over-smoothness caused by a larger $\lambda$.
3.3 Semantic Enhancement Score Distillation
@@
Current text-to-3D generation methods like DreamFusion [37] often optimize a 3D representation parameterized by $\theta$ so that rendered images $\mathbf{x}=g(\theta)$ resemble 2D samples produced by a pre-trained T2I diffusion model for a given text prompt $y$. $g$ functions as a differentiable renderer. The T2I diffusion model $\phi$ predicts the sampled noise $\epsilon_{\phi}(\mathbf{x}_{t};y,t)$ of a rendered image $\mathbf{x}_{t}$ at a noise level $t$ for the text input $y$. To make rendered images follow the text-conditioned distribution in Stable Diffusion, the SDS loss updates $\theta$ by estimating the gradient:
@@
However, the SDS loss cannot effectively supervise sub-object structure generation due to the issue of semantic coupling in full objects. For example, when using the SDS loss to generate a tortoise shell, it also usually generates the tortoise’s tail and head, causing semantic coupling (Figure 7). We believe that the issue of semantic coupling in SDS stems from the training data of Stable Diffusion, in which most images contain full objects rather than separate parts. Therefore, the semantics of full objects often appear in the target distribution conditioned on text only describing sub-object structures.
@@
A straightforward approach is using negative distributions via Classifier Score Distillation (CSD) [59] or Variational Score Distillation (VSD) [54] to suppress coupled semantics. CSD employs predefined negative prompts, resulting in more accurate negative distributions than those adaptively learned by VSD [59]:
@@
$\nabla_{\theta}\mathcal{L}_{\text{CSD}}(\phi,\mathbf{x})$ $=\mathbb{E}_{t,\epsilon,c}[(\omega_{\text{pos}}\cdot\epsilon_{\phi}(\mathbf{x}_{t};y,t)$
@@
$-\omega_{\text{neg}}\cdot\epsilon_{\phi}(\mathbf{x}_{t};y_{\text{neg}},t))\frac{\partial\mathbf{x}}{\partial\theta}],$ (5)
where $\omega_{\text{pos}}$ and $\omega_{\text{neg}}$ denote different weights for positive and negative prompts. It requires two separate inferences with positive and negative prompts to obtain two distributions, which are then subtracted to suppress coupled semantics. However, our experiments show that CSD is ineffective in decoupling semantics because the negative prompt cannot accurately represent the undesirable coupled semantics in the positive prompt (Figure 7). This results in noisy guidance, making CSD less effective in decoupling semantics.
Unlike semantic suppression in CSD, we propose a semantic enhancement method to mitigate semantic coupling by enhancing the semantics of part-related words. This can lead to a more accurate and stable target distribution, as shown in Section 4.3. Our key design is to apply weighted blending to the tokens in the original prompt to obtain a semantically focused text embedding, which serves as stable guidance for the optimization process. We define the Semantic Enhancement SDS loss as:
@@
$\small\nabla_{\theta}\mathcal{L}_{\text{SE}}(\phi,\mathbf{x})=\mathbb{E}_{t,\epsilon,c}\left[\omega(t)\left(\epsilon^{*}_{\phi}(\mathbf{x}_{t};y,t)-\epsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right],$ (6)
@@
where $\epsilon^{*}_{\phi}(\cdot)$ uses a text embedding weighted by Compel [48]. Specifically, we assign a weight $s$ to each word in the prompt and compute the weighted embedding $e_{\text{w}}$ for each word by blending the original word embedding $e$ and the empty text embedding $e_{0}$: $e_{\text{w}}=e_{0}+s\cdot(e-e_{0})$. By concatenating the weighted embedding of each word in sequence, we obtain the final semantically focused text embedding. In our experiments, we found that using $s={1.1}^{2}$ for words that require semantic enhancement can achieve stable optimization and effectively alleviate the issue of semantic coupling.
@@
Figure 5: Qualitative comparisons of generated brushes for surface details. Our method captures geometry details guided by text, effectively preserving the surface structure and avoiding mesh distortion.